# Complexity Scaling Laws for Neural Models using Combinatorial Optimization

### v1

@@ H.2 BOHB results
Figure 18 summarizes the BOHB search results used to inform our HP selection. Using larger model widths, a tight range near $10^{-4}$ for max learning rate, and a small gradient norm clip less than 1 are most important for achieving high performance with fewer model updates. Convergence was particularly sensitive to gradient norm clip values greater than 1. In preliminary attempts without gradient clipping, pruning simply selected trials which most recently diverged.
@@
Optuna tracks objective scores by the final value achieved at the maximum update budget, not the best value achieved over the trial, which poses a question of statistical significance in near-optimal improvements. For example, trial 300 had the best outcome with a suboptimality gap of 0.169 at completion. But its minimum validation score over training was slightly below this at 0.166, though this happens to be the minimum score reported over all validations for all trials, which is encouraging. We evaluated the impact of our validation dataset size relative to these small performance deltas. This was performed retrospectively using the 50-node model from our main experiments since no checkpoints were saved during HPO. We rescored with 100 randomly sampled datasets of 10,000 problems each, obtaining a standard deviation of less than 0.003 suboptimality. The difference in suboptimality between the top two trials is over four standard deviations. Even comparing the final validation score of the top trial with the minimum score of the second-place trial returns a difference of over two standard deviations, demonstrating our sample size is statistically meaningful assuming validation scores are normally distributed.
@@ Appendix I Proof of scaling relations and limits
Note that each $\lVert X\rVert_{2}^{(p_{k},q_{k})}$ variable shares sampled coordinates with their adjacent edges in the sum, meaning $\lVert X\rVert_{2}^{T}$ is not a sum of independent variables. But these random edge norms are still identically distributed given the choice of starting node $p_{1}$ is arbitrary. Thus,
@@
which slightly complicates this proof. However, we can make use of the fact that dependencies are limited to adjacent edge variables. We start with the following equation:
@@
Given covariances between adjacent edge lengths are identically distributed, we obtain
@@
where $(p,q,s)$ denotes an arbitrary 3-node sequence in the tour with edges $(p,q)$ and $(q,s)$. ∎
@@
[…]
@@
[…]
@@
[…]
@@
since the product inside the expected value is identically distributed for each summation term over $d$. Condensing, we arrive at
@@
[…]
@@
[…]
@@
Here we leverage an existing derivation that bounds the limit of expected absolute contrast for a fixed number of nodes. Taking the norm from an arbitrary query point to each node coordinate, absolute contrast is the maximum difference between norms. Specifying the Euclidean case for Corollary 2 in Aggarwal et al. [27], the bound can be formally described as follows:
@@
where node $p$ is our chosen query point, [^13] Aggarwal et al. consistently use the origin as their query point but without loss of generality. Our upper bound term is scaled by $(n-2)$ to account for removing node $p$ from the end point set $Q$. $Q$ is the set of remaining nodes such that $p\notin Q$, and $\lambda$ is a non-negative constant. This bound applies to any node coordinate distribution satisfying i.i.d. sampling for each dimension.
@@
00 0 $\leq\lVert X\rVert_{2}^{T}-\lVert X\rVert_{2}^{T^{*}}$
00 0 $\leq\lim_{d\to\infty}\operatorname{\mathbb{E}}\left[\lVert X\rVert_{2}^{T}-\lVert X\rVert_{2}^{T^{*}}\right]$

### v2

@@ H.2 BOHB results
Figure 18 summarizes the BOHB search results that were used to inform our HP selection. Using larger model widths, a tight range near $10^{-4}$ for max learning rate, and a small gradient norm clip less than 1 are most important for achieving high performance with fewer model updates. Convergence was particularly sensitive to gradient norm clip values greater than 1. In preliminary attempts without gradient clipping, pruning simply selected trials which most recently diverged.
@@
Optuna tracks objective scores by the final value achieved at the maximum update budget, not the best value achieved over the trial, which poses a question of statistical significance in near-optimal improvements. For example, trial 300 had the best outcome with a suboptimality gap of 0.169 at completion. But its minimum validation score over training was slightly below this at 0.166, though this happens to be the minimum score reported over all validations for all trials, which is encouraging. We evaluated the impact of our validation dataset size relative to these small performance deltas. This evaluation was performed retrospectively using the 50-node model from our main experiments since no checkpoints were saved during HPO. We rescored with 100 randomly sampled datasets of 10,000 problems each, obtaining a standard deviation of less than 0.003 suboptimality. The difference in suboptimality between the top two trials is over four standard deviations. Even comparing the final validation score of the top trial with the minimum score of the second-place trial returns a difference of over two standard deviations, demonstrating that our sample size is statistically meaningful assuming validation scores are normally distributed.
@@ Appendix I Proof of scaling relations and limits
Note that each $\lVert X\rVert_{2}^{(p_{k},q_{k})}$ variable shares sampled coordinates with their adjacent edges in the sum, meaning $\lVert X\rVert_{2}^{T}$ is not a sum of independent variables. But these random edge norms are still identically distributed given that the choice of starting node $p_{1}$ is arbitrary. Thus,
@@
This slightly complicates the proof. However, we can make use of the fact that dependencies are limited to adjacent edge variables. We start with the following equation:
@@
Given that covariances between adjacent edge lengths are identically distributed, we obtain
@@
where $(p,q,s)$ denotes an arbitrary 3-node sequence in the tour with edges $(p,q)$ and $(q,s)$, and $\left(\xi^{(p,q,s)}\right)^{2}$ is defined as the summation term obtained above. ∎
@@
$\mu_{{\chi^{2}}}=\operatorname{\mathbb{E}}\left[\left(\chi^{(p,q)}\right)^{2}\right]=\int_{0}^{1}\int_{0}^{1}f(\omega)f(\rho)(\rho-\omega)^{2}\,d\rho\,d\omega$
@@
$\mu_{{\chi^{2}}}=\int_{0}^{1}\int_{0}^{1}(\rho-\omega)^{2}\,d\rho\,d\omega=\int_{0}^{1}\left(\omega^{2}-\omega+\frac{1}{3}\right)d\omega=\frac{1}{6}$
@@
Extending existing analysis, it is simple to show that the previous upper bound becomes arbitrarily tight in the limit as $d\to\infty$. Specifying the Euclidean norm case121212p=2p=2 in the cited work. Our usage of $p$ as the norm starting node is unrelated. for the proof of Lemma 1 in François et al. [28] (Section 5.1.1), it directly follows that
@@
since the product inside the expected value is identically distributed for each summation term over $d$. We define this summation term as $\left(\varphi^{(p,q,s)}\right)^{2}$. Condensing, we arrive at
@@
$=\frac{\lim\limits_{d\to\infty}\frac{\operatorname{\mathbb{E}}\left[\left(\sum_{k=1}^{d}\left(\left(\chi_{k}^{(p,q)}\right)^{2}-\mu_{{\chi^{2}}}\right)\right)\left(\sum_{j=1}^{d}\left(\left(\chi_{j}^{(q,s)}\right)^{2}-\mu_{{\chi^{2}}}\right)\right)\right]}{d}}{4\,\mu_{{\chi^{2}}}}\quad\text{via Fatou–Lebesgue theorem}$
@@
$\left(\sigma_{{\chi^{2}}}\right)^{2}=$ $\int_{0}^{1}\int_{0}^{1}f(\omega)f(\rho)((\rho-\omega)^{2}-\mu_{{\chi^{2}}})^{2}d\rho\,d\omega=\int_{0}^{1}\int_{0}^{1}\left((\rho-\omega)^{2}-\frac{1}{6}\right)^{2}d\rho\,d\omega=\frac{7}{180}$
$\left(\varphi^{(p,q,s)}\right)^{2}$ $=\int_{0}^{1}\int_{0}^{1}\int_{0}^{1}f(\tau)f(\omega)f(\rho)((\rho-\omega)^{2}-\mu_{{\chi^{2}}})((\omega-\tau)^{2}-\mu_{{\chi^{2}}})d\rho\,d\omega\,d\tau$
$=\int_{0}^{1}\int_{0}^{1}\int_{0}^{1}\left((\rho-\omega)^{2}-\frac{1}{6}\right)\left((\omega-\tau)^{2}-\frac{1}{6}\right)d\rho\,d\omega\,d\tau$
@@
Here we leverage an existing derivation that bounds the limit of expected absolute contrast for a fixed number of nodes. Taking the norm from an arbitrary query point to each node coordinate, absolute contrast is defined as the maximum difference between norms. Specifying the Euclidean case for Corollary 2 in Aggarwal et al. [27], the bound can be formally described as follows:
@@
where node $p$ is our chosen query point, [^14] Aggarwal et al. consistently use the origin as their query point but without loss of generality. Our upper bound term is scaled by (n2)(n-2) to account for removing node $p$ from the end point set $Q$. $Q$ is the set of remaining nodes such that $p\notin Q$, and $\lambda$ is a non-negative constant. This bound applies to any node coordinate distribution satisfying i.i.d. sampling for each dimension.
@@
$0$ $\leq\lVert X\rVert_{2}^{T}-\lVert X\rVert_{2}^{T^{*}}$
$0$ $\leq\lim_{d\to\infty}\operatorname{\mathbb{E}}\left[\lVert X\rVert_{2}^{T}-\lVert X\rVert_{2}^{T^{*}}\right]$